OpenStudy (anonymous):

How would I simplify the square root 98 times x to the fifth power.

OpenStudy (anonymous):

umm all u gotta do is find \[\sqrt{98}\]

OpenStudy (anonymous):

which is 2times 49 and 49 equals 7 times 7 so the equation is 7\[\sqrt{2}\]

OpenStudy (anonymous):

\[7\]

OpenStudy (anonymous):

\[7\sqrt{2}\]

OpenStudy (anonymous):

so 7x^5 radical 2 is answer

OpenStudy (anonymous):

I have 7x^2 radical 2 as a choice, should I do that one?:o

OpenStudy (anonymous):

how because \[7\sqrt{2}timesx ^{5}is basically 7x ^{5}\]

OpenStudy (anonymous):

radical 2

OpenStudy (anonymous):

I don't know, it's just the closest answer choice to what you said.

OpenStudy (anonymous):

i mean when u multiplay x^5 and radical 98 u just find radical 98 which is 7 radical 2 and u just add x^5 to 7 which is 7x^5radical 2

OpenStudy (anonymous):

when its 2x^5 times 7 than itll be 14x^5

OpenStudy (anonymous):

as in example

OpenStudy (anonymous):

Alright, thanks!

OpenStudy (anonymous):

np